OpenStudy (anonymous):

what is the inverse of f(x)=8-x^2

OpenStudy (anonymous):

say y=8-x^2 and try writing x in terms of y

OpenStudy (anonymous):

construct and simplify the difference quotient f(h+h)-f(x)/h The difference quotient is= f(x)=x^2-8 Now, we can just substitute these into the expression: (f(x+h)-f(x))/h=(%28%28x%2Bh%29%5E2-8%29-%28x%5E2-8%29%29%2Fh+)/h=%28x%5E2%2B2xh%2Bh%5E2-8-x%5E2%2B8%29%2Fh=%282xh%2Bh%5E2%29%2Fh=2x%2Bh That is the answer. In calculus, we would take the limit as h goes to zero. If h is 0, 2x+h=2x. This gives the derivative of the function f(x)=x^2+c, c arbitrary. I hope you find that information interesting

OpenStudy (anonymous):

momo1999, you're finding the derivative. they only need the inverse. no calculus involved.

OpenStudy (anonymous):

f (x) = 2x – 1 y = 2x – 1 y + 1 = 2x (y + 1)/2 = x (x + 1)/2 = y (x + 1)/2 = f –1(x) like that